40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Wed Mar 14 05:26:39 2012, interferomter tack force, Update, General, daytime tasks 
    Reply  Wed Mar 14 14:09:01 2012, interferomter tack force, Update, General, daytime tasks 
    Reply  Wed Mar 14 16:39:02 2012, Suresh, Update, General, REFL165 signal was not reaching demod board : Fixed Test_CDS_Calibration.png
Message ID: 6416     Entry time: Wed Mar 14 14:09:01 2012     In reply to: 6412
Author: interferomter tack force 
Type: Update 
Category: General 
Subject: daytime tasks 

Quote:

The following tasks need to be done in the daytime tomorrow.

  • Hook up the DC output of the Y green BBPD on the PSL table to an ADC channel (Jamie / Steve)
  • Install fancy suspension matrices on PRM and ITMX [#6365] (Jenne)
  • Check if the REFL165 RFPD is healthy or not (Suresh / Koji)
    • According to a simulation the REFL165 demod signal should show similar amount of the signal to that of REFL33.
    • But right now it is showing super tiny signals [#6403]

 For ITMX, I used the values from the conlog:

2011/08/12,20:10:12 utc 'C1:SUS[-_]ITMX[-_]INMATRIX'
These are the latest values in the conlog that aren't the basic matricies.  Even though we did a round of diagonalization in Sept, and the 
matricies are saved in a .mat file, it doesn't look like we used the ITMX matrix from that time.

For PRM, I used the matricies that were saved in InputMatricies_16Sept2011.mat, in the peakFit folder, since I couldn't find anything in the Conlog other than the basic matricies.

 

UPDATE:  I didn't actually count the number of oscillations until the optics were damped, so I don't have an actual number for the Q, but I feel good about the damping, after having kicked POS of both ITMX and PRM and watching the sensors.

ELOG V3.1.3-